please help me i am stuck on this question.......
if r^8=5 and r^7=3/t^2,what is the value of r in terms of t?

Answer by rothauserc(4718)

r^7 = 3/t^2
r^8 = 5
r^7 * r = 5
(3/t^2) * r = 5
t^2 is denominator then multiply both sides of = by t^2
3r = 5t^2
r = 5t^2/3